x^x + y^y = 3
x + y = 5

Find the value of x and y
1
Expert's answer
2020-11-27T10:16:00-0500

The question is actually

"x^x + y^y = 31\\

x + y = 5"

not,


"x^x + y^y = 3\\

x + y = 5\\"



The best way to solve this is graphically,

Table "x^x + y^y = 31\\"

"x|y\\
0|3.04\\
1|3.04\\
2|3\\
3|0\\"


Table "x + y = 5"


"x|y\\
0|5\\
1|4\\
2|3\\
3|2\\"




The points of intersection are (2,3) and (3,2)

"\therefore" "x=2" or "3"

"y=3" or "2"